Circular No. 8827

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 E3 (LINEAR) E. J. Christensen reports his recovery of comet P/1999 J5 (cf. IAUC

7201) on images obtained in the course of the Mount Lemmon Survey on Mar. 9 by himself and on Mar. 10 by R. A. Kowalski. The images were not obviously cometary, although they may have been slightly elongated toward p.a. 235^{o} on the first night. On the second night, 30-second integrations in good seeing revealed neither coma nor tail. G. V. Williams, Minor Planet Center, identified further observations in Mt. Lemmon incidental asteroidal astrometry by Kowalski on Apr. 7.

2007 UT  α_{2000}  δ_{2000}  Mag.
Mar. 9.37740  11^{h}18^{m}40.^{s}14  +19^{o}55′38′′.2  21.2
9.38374  11 18 39.96  +19 55 40.8  21.1
9.39006  11 18 39.74  +19 55 42.6  21.1
9.39652  11 18 39.49  +19 55 44.2  21.0
10.31285  11 18 07.57  +20 00 05.7
10.33449  11 18 06.78  +20 00 11.9
10.34527  11 18 06.39  +20 00 15.3
Apr. 7.26289  11 03 34.84  +21 26 40.2  22.0
7.26917  11 03 34.66  +21 26 39.1  21.2
7.27556  11 03 34.48  +21 26 40.5  20.9
7.28187  11 03 34.34  +21 26 40.4  20.8

The indicated correction to the prediction on MPC 54170 is ΔT = −0.8 day. The following improved orbital elements have been computed by Williams from 72 observations, 1999–2007:

Epoch = 1999 May 22.0 TT

T = 1999 May 12.1310 TT  ω = 132.^{o}1799
e = 0.169470  Ω = 112.0187 }2000.0
q = 3.712711 AU  i = 13.7164

a = 4.470289 AU n^{o} = 0.1042799 P = 9.45 years Epoch = 2008 Oct. 21.0 TT

T = 2008 Oct. 6.3021 TT  ω = 131.^{o}9564
e = 0.170548  Ω = 112.0027 }2000.0
q = 3.693255 AU  i = 13.7323

a = 4.452644 AU n^{o} = 0.1049004 P = 9.40 years

2007 April 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green